OpenStudy (anonymous):

how many minutes does it take a motorist traveling at 40 miles an hour to reach a point 2/5 of a mile away @zepdrix @ganeshie8 @nincompoop A) 4/15 B) 2/3 C) 1 D) 4 E) 3/5

OpenStudy (anonymous):

rate problem...

OpenStudy (anonymous):

Distance = rate x time

OpenStudy (anonymous):

@Orion1213 Distance= rate* time is equal to 16=40* 2/5 which is not an answer choice

OpenStudy (anonymous):

your looking for time, you have to re-arrange the formula i gave...

OpenStudy (anonymous):

distance = 2/5 of a mile... and rate = 40 miles an hour = 40 miles per hour... so what will you do to put your answer in time unit?

OpenStudy (anonymous):

D/ R= T ?

OpenStudy (anonymous):

yup that's correct...

OpenStudy (anonymous):

But that equals 1/100 which is not an answer choice

OpenStudy (anonymous):

take note the unit of your time, it's not yet in minutes... how many minutes do we have in an hour?

OpenStudy (anonymous):

60 minutes

OpenStudy (anonymous):

and so... 1/100 hour = in minutes?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

no... it is 60/100 minutes... simplify it further...